x is a random variable with with a positive mean and for which E[x(x+1)]=80 and Var(x)=50. what is E[x]
Answer #1

Solution:

E[x(x+1)] = 80

E[x2 + x ] = 80

E[x2 ] + E[x] = 80

E[x2 ] = 80 - E[x]

Now, we know that

Var(x) = E[x2 ] - [E(x)]2

50 = 80 - E[x]  - [E(x)]2  

[E(x)]2 + E[x ] -30 = 0

The roots of this quadratic equations are 6 and -5

But it is given that the mean of x is positive.i.e. E[x ] is positive.

  E[x ] = 6